OpenStudy (anonymous):

if two adjacent anges form a straight line, and the ratio of the larger angle to the smaller is 7 to 3, what is the measure in degrees of the smaller angle? (1)72 (2)54 (3)18 (4)126

OpenStudy (anonymous):

*ANGLES

OpenStudy (mertsj):

x+y = 180 x/y=7/3

OpenStudy (anonymous):

okay

OpenStudy (mertsj):

Can you solve it from there?

OpenStudy (anonymous):

I'll see

OpenStudy (anonymous):

No, I don't know :(

OpenStudy (mertsj):